OpenStudy (anonymous):

Find the differential of y=sqrt(9+t^2)

OpenStudy (accessdenied):

y = sqrt(9 + t^2) = (9 + t^2)^(1/2) putting it into exponential form to use product rule We could use chain rule, where f(x) = x^(1/2) and g(x) = 9 + t^2 y' = f'(g(x)) * g'(x) f'(x) = (1/2)x^(-1/2) = (1/2x^(1/2)) So, we plug in g(x) for x f'(g(x)) = (1/2(9 + t^2)^(1/2)) g'(x) = 2t So, f'(g(x)) * g'(x) = 2t * (1/2(9 + t^2)^(1/2)) = t/(9 + t^2)^(1/2) \[ y' = \frac{t}{\sqrt{9 + t^{2}}} \]

OpenStudy (accessdenied):

sorry, not 'product rule', i mean 'power rule'.

OpenStudy (anonymous):

Thank you AccessDenied! :D

OpenStudy (accessdenied):

No problem. Glad to help! :)